OpenStudy (anonymous):

Solve: x^2-81=0 Please show how to solve step by step.

OpenStudy (anonymous):

\[x^2=81\] \[x=9\] or \[x=-9\]

OpenStudy (anonymous):

Factor (x+9)(x-9)=0 solve: (x+9)=0 x=-9 (x-9)=0 x=9